OpenStudy (anonymous):

can anyone find out the 8th and the nth term of the sequence "3,6,7,-2" and the nth term of the sequence "7,5,3,1..." ??

OpenStudy (anonymous):

for the second one t(n) = 7 - 2(n-1)

OpenStudy (anonymous):

thanks :) and do you have any idea about the other one? it seems to be very vaque :P

OpenStudy (anonymous):

the first one i cant figure out

OpenStudy (amistre64):

3 3+ 3 = 6 6+ 1 = 7 7+(-9) = -2 1(3) + (a) = 6 2(3) + (a) = 7 3(3) + (a) = -2 3x + (a) = y 0,3 1,6 2,7 3,-2 im wondering if we match a curve to these points if that would generate a function ....

OpenStudy (amistre64):

well, actually: "3,6,7,-2" only has 4 terms to it; without the ... there is no other terms to find

OpenStudy (anonymous):

hahahahaha

OpenStudy (amistre64):

:) unless it was a typo, thats the best answer I can come up with ;)

OpenStudy (anonymous):

actually, it had to do with two other sequences. there was 4,9,16,25 and 1,3,9,27. so this sequence is the first one minus the second one. 4-3, 9-3, 16-9, 25-27, etc.